A 2.5 g sample of KClO3 was heated. The liberated oxygen gas was collected over water at 24 degree Celsius and 755 millimeter mercury. What was the mass and volume of the oxygen gas collected in the set up?


1
Expert's answer
2021-01-29T08:21:03-0500
"P1V1\/T1=P0V0\/T0"

755 mm Hg = 100.658 kPa


"V1=P0V0T1\/(T0*P1)""2KClO3\to2KCl+3O2"

n(KClO3)=2.5/122.5=0.02 mol

n(O2)=3*0.02/2=0.03 mol

m(O2)=0.03*32=0.96 g

V(O2)=0.03*22.4=0.672 L


"VO2=101.325*0.672*297\/(273*100.658)=0.736 L"